Bramble Corp. purchased machinery that was installed and ready for use on January 3, 2017, at a total cost of $242000. Salvage value was estimated at $28000. The machinery will be depreciated over five years using the double-declining balance method. For the year 2018, Bramble should record depreciation expense on this machinery of
Answer #1
Depreciation rate under double decline method
depreciation rate = 1/estimated life *2
depreciation rate = 1/5years*2
Depreciation rate = 40%
Depreciation expenses 2017= 242000*40%
Depreciation expenses 2017 = 96800
Depreciation expenses 2018= (cost of assets-depreciation exp 2017)*40%
Depreciation expenses 2018 = (242000-96800)*40%
Depreciation expenses 2018 = 58080